# Predicting the Curie temperature in substitutionally disordered alloys using a first-principles based model

### v1

@@ 5 Discussion
For the Fe1-xCox system, our model shows especially good agreement in the cobalt-rich fcc region, where the deviation is at most 60 K. In the bcc region, the error stays below $80$ K. Overall, the errors of the Fe-containing binary alloys are consistently within our previous estimate of the mean absolute error, 126 K, of the model [20] and the trends with the alloy composition are well described.
@@
Our model successfully captures the concentration dependence of $T_{\text{C}}$ in NixCu1-xMnSb, aligning with experimental observations. The error is generally within 150 K, with a maximum of approximately 170 K. This is comparable to the effects of 5% intersite disorder in this half-Heusler alloy, as reported in Ref. [5], but not considered in the present work. The predicted $T_{\text{C}}$ values display a nearly linear decrease as copper content increases especially in the copper-rich region, with a slight underestimation compared to experimental data. For the NixMnSb system, the predictions are very close to the experimental values, with the largest error being around 100 K at $x=1.75$. This is also in the composition range at which Ref. [56] observed increasing disorder form the C1b structure.
In Fig. 1, a hint of an increase in $T_{\text{C}}$ as $x$ increases is seen for the Ti1-xCrxN system. The incline is comparable to the experimental incline, however, with a significant overestimation of the absolute value of $T_{\text{C}}$. The magnetic moments of the chromium atoms did not align completely ferromagnetically in the ground-state search; the energies are, however, relatively similar to the ferromagnetic case. Consistently over the different $x$-values, the ground state resulted in a slightly lower energy compared to a perfect ferromagnetic configuration. The largest difference is for $x=0.5$, which is also the upper limit for ferromagnetism in this system [57, 53, 6]. These differences do not change the predicted $T_{\text{C}}$ significantly.
For Ti1-xCrxN the experimental Curie temperatures are relatively low. One should note that the model of Eq. (1) has a lower boundary of 124 K, since $\Delta E$ is never below zero. Approaching a Curie temperature near 124 K will therefore naturally lead to problems in our predictions. The experimental $T_{\text{C}}$ in Ti1-xCrxN is at most 140 K, which is very close to the lower limit of our model. This overestimation is therefore likely due to the inherent limitations of the model. A way of resolving this could be by introducing a correction term enabling differentiation between the low and high $T_{\text{C}}$ regions. The design of our model is, however, based on the essential physical assumptions described briefly in Sec. 2.2 and in more detail in Ref. [20]. Introducing a higher level of complexity into the model structure in an attempt to resolve this issue would potentially obscure the physics assumptions on which this model is based.
@@
As the technetium content increases, $T_{\text{C}}$ decreases in bcc Fe1-xTcx, as shown in Fig. 4. This trend corresponds well to the behavior observed in experimentally studied systems such as Fe1-xRux and Fe1-xMox, which also contain $4d$ elements that lower the Curie temperature [59, 60]. However, unlike technetium, these elements are not radioactive. Molybdenum, in particular, has a very similar electronic structure to technetium, with both elements having completely half-filled $4d$ orbitals, which suggests that technetium might contribute to the magnetic interactions in iron alloys in a way comparable to molybdenum. Similarly, Fe1-xRex exhibits a comparable electronic structure, but with $5d$ electrons. In this case, a reduction in the Curie temperature is also observed with increasing rhenium content [61].
6 Conclusions
In this work we have applied the model developed in our previous work Ref. [20] to the substitutional disordered alloys bcc and fcc Fe1-xCox, bcc Fe1-xCrx, bcc Fe1-xVx, NixCu1-xMnSb, NixMnSb, Ti1-xCrxN, fcc Co1-xAlx and the experimentally unexplored bcc Fe1-xTcx. The model accurately predicts the trends of the variation of $T_{\text{C}}$ with composition and in some cases is in good absolute agreement with experimental values. Our model has a lower limit of 124 K which naturally becomes apparent for systems with low $T_{\text{C}}$ such as Ti1-xCrxN. However, the model still predicts an accurate $\partial T_{\text{C}}/\partial x$ in this system. Systems with a more itinerant type of magnetism, e.g., Co1-xAlx, are not well described by this method, as it depends mainly on the rotational disorder of robust magnetic moments to describe the paramagnetic state. Overall, the approach provides a valuable and computationally efficient tool for predicting trends in $T_{\text{C}}$, offering insights that can guide experimental studies and alloy design. This approach not only increases computational efficiency compared to other theoretical methods but also allows for a broader applicability, without the need for extensive system-specific human intervention.

### v2

@@ V Discussion
For the Fe1-xCox system, our model shows especially good agreement in the cobalt-rich fcc region, where the deviation is at most 60 K. In the bcc region, the error stays below $80$ K. Overall, the errors of the Fe-containing binary alloys are consistently within our previous estimate of the mean absolute error, 126 K, of the model brannvall_predicting_2024 and the trends with the alloy composition are well described.
@@
Our model successfully captures the concentration dependence of $T_{\text{C}}$ in NixCu1-xMnSb, aligning with experimental observations. The error is generally within 150 K, with a maximum of approximately 170 K. This is comparable to the effects of 5% intersite disorder in this half-Heusler alloy, as reported in Ref. alling_effect_2009 , but not considered in the present work. The predicted $T_{\text{C}}$ values display a nearly linear decrease as copper content increases especially in the copper-rich region, with a slight underestimation compared to experimental data. For the NixMnSb system, the predictions are very close to the experimental values, with the largest error being around 100 K at $x=1.75$. This is also in the composition range at which Ref. webster_chemical_1984 observed increasing disorder form the C1b structure.
In Fig. 1, a hint of an increase in $T_{\text{C}}$ as $x$ increases is seen for the Ti1-xCrxN system. The incline is comparable to the experimental incline, however, with a significant overestimation of the absolute value of $T_{\text{C}}$. The magnetic moments of the chromium atoms did not align completely ferromagnetically in the ground-state search; the energies are, however, relatively similar to the ferromagnetic case. Consistently over the different $x$-values, the ground state resulted in a slightly lower energy compared to a perfect ferromagnetic configuration. The largest difference is for $x=0.5$, which is also the upper limit for ferromagnetism in this system aivazov_magnetic_1975 ; inumaru_ferromagnetic_2007 ; alling_theory_2010 . These differences do not change the predicted $T_{\text{C}}$ significantly.
For Ti1-xCrxN the experimental Curie temperatures are relatively low. One should note that the model of Eq. (1) has a lower boundary of 124 K, since $\Delta E$ is never below zero. Approaching a Curie temperature near 124 K will therefore naturally lead to problems in our predictions. The experimental $T_{\text{C}}$ in Ti1-xCrxN is at most 140 K, which is very close to the lower limit of our model. This overestimation is therefore likely due to the inherent limitations of the model. A way of resolving this could be by introducing a correction term enabling differentiation between the low and high $T_{\text{C}}$ regions. The design of our model is, however, based on the essential physical assumptions described briefly in Sec. II.2 and in more detail in Ref. brannvall_predicting_2024 . Introducing a higher level of complexity into the model structure in an attempt to resolve this issue would potentially obscure the physics assumptions on which this model is based.
@@
As the technetium content increases, $T_{\text{C}}$ decreases in bcc Fe1-xTcx, as shown in Fig. 4. This trend corresponds well to the behavior observed in experimentally studied systems such as Fe1-xRux and Fe1-xMox, which also contain $4d$ elements that lower the Curie temperature pottker_mossbauer_2004 ; arajs_ferromagnetic_1965 . However, unlike technetium, these elements are not radioactive. Molybdenum, in particular, has a very similar electronic structure to technetium, with both elements having completely half-filled $4d$ orbitals, which suggests that technetium might contribute to the magnetic interactions in iron alloys in a way comparable to molybdenum. Similarly, Fe1-xRex exhibits a comparable electronic structure, but with $5d$ electrons. In this case, a reduction in the Curie temperature is also observed with increasing rhenium content cieslak_discovery_2014 .
VI Conclusions
In this work we have applied the model developed in our previous work Ref. brannvall_predicting_2024 to the substitutional disordered alloys bcc and fcc Fe1-xCox, bcc Fe1-xCrx, bcc Fe1-xVx, NixCu1-xMnSb, NixMnSb, Ti1-xCrxN, fcc Co1-xAlx and the experimentally unexplored bcc Fe1-xTcx. The model accurately predicts the trends of the variation of $T_{\text{C}}$ with composition and in some cases is in good absolute agreement with experimental values. Our model has a lower limit of 124 K which naturally becomes apparent for systems with low $T_{\text{C}}$ such as Ti1-xCrxN. However, the model still predicts an accurate $\partial T_{\text{C}}/\partial x$ in this system. Systems with a more itinerant type of magnetism, e.g., Co1-xAlx, are not well described by this method, as it depends mainly on the rotational disorder of robust magnetic moments to describe the paramagnetic state. Overall, the approach provides a valuable and computationally efficient tool for predicting trends in $T_{\text{C}}$, offering insights that can guide experimental studies and alloy design. This approach not only increases computational efficiency compared to other theoretical methods but also allows for a broader applicability, without the need for extensive system-specific human intervention.
@@ Appendix A Numerical values of predicted and
Data availability
Python scripts used to extract the data from density functional theory calculations are openly available in the DFT_based_CurieTemp_predictions repository at https://github.com/marbr639/DFT_based_CurieTemp
_predictions
. Other data that support the findings of this study are available upon request.